OpenStudy (anonymous):

Find the surface area of the prism. A. 2299 mm2 B. 2954.5 mm2 C. 3173 mm2 D. 498,180 mm2

OpenStudy (anonymous):

http://prntscr.com/9q51r3

OpenStudy (anonymous):

@purple_pink

OpenStudy (anonymous):

@ParthKohli

OpenStudy (purple_pink):

lemme calculate

OpenStudy (studygurl14):

SURFACE AREA = \(\Large 2(B) +Ph\) B = area of base (in this case, the triangle) P = perimeter of base h = hieght

OpenStudy (anonymous):

ol

OpenStudy (anonymous):

ok i mean

OpenStudy (studygurl14):

:) So...plug in known values... \(\Large SA = 2(\frac{19(23)}{2})+(23+19+30)(38)\)

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

ok no what

OpenStudy (anonymous):

purple_pink are u still here

OpenStudy (purple_pink):

ya hold on almost done

OpenStudy (anonymous):

oh ok

OpenStudy (studygurl14):

I laid it out for you. Just plug it all into a calulator if you don't want to do it yourself

OpenStudy (anonymous):

oh ok ty

OpenStudy (purple_pink):

ok you still here?

OpenStudy (purple_pink):

@littleelfie ??

OpenStudy (anonymous):

yes

OpenStudy (purple_pink):

your answer is C Because, 2(1/2)(23)(19) + (38)(19) + (38)(30) + (38)(23) = 3173 3173 is C Medal?

OpenStudy (anonymous):

there you go

OpenStudy (purple_pink):

thanks

OpenStudy (anonymous):

np